OpenStudy (anonymous):

Help! Please explain! Find the roots of the polynomial equation. x2 – 14x – 32 = 0 A. 2 and –16 B. –6 and 6 C. –2 and 16 D. –9 and 8

OpenStudy (***[isuru]***):

hi, factorize the expression first.. x2 – 14x – 32 = 0 x^2 - 16x + 2x -32 =0 x( x - 16 ) +2 ( x -16) = 0 (x + 2)( x -16) = 0 did u get it so far ??

OpenStudy (anonymous):

yes ok

OpenStudy (***[isuru]***):

so... now we got (x + 2)( x -16) = 0 which means either (x + 2) = 0 or ( x -16) = 0 from that we can say... x = -2 or x = 16 I think now u have ur answers lol

OpenStudy (anonymous):

ok thnx